# TransformerFAM: Feedback attention is working memory

### v2

@@ 3.2 PassKey Retrieval
The PassKey retrieval task is a recent benchmark used in several long-context transformer papers (Mohtashami & Jaggi, 2023; Tworkowski et al., 2023; Chen et al., 2023). In this task, a passkey is presented at the beginning, followed by a very long filler context. Finally, a question about the passkey is asked, as shown in Figure 11 in Section D.1.
@@
As shown in Figure 3, TransformerFAM was able to perfectly solve the task with a filler context of up to 260k tokens. In the figure, MX denotes the number of BSWA memory segments. The performance of TransformerBSWA improves significantly up to M2, after which it saturates. The performance of M12 also drops significantly after 20k tokens. The theoretical receptive field of M2 is 36k (i.e. depth(18) $\times$ memory segment(2k)), but the effective receptive field is much shorter.
Figure 3: PassKey Retrieval: Performance across different Transformer models and memory segment configurations. MX denotes the number of BSWA memory segments. FAM represents TransformerFAM with 0 memory segments. TransformerFAM successfully solves the task.
In Figure 3, it is important to compare M1_SG and M1. M1_SG has a stop gradient applied to one memory segment, which limits the receptive field to the window size. This is because the model cannot learn which contextual information stored in the memory segment will have a good result later. SWA with back propagation through time (BPTT) functions similarly to a time-limited RNN. It is common to use stop gradients on the memory segments of TransformerXL (Dai et al., 2019; Chevalier et al., 2023). However, we recommend against this practice.
In Section B.4.6, we compare our work with recent Transformer with memory papers (Bulatov et al., 2022; Chevalier et al., 2023).
@@ 3.3 Long Context Tasks
Gemini (Team et al., 2023) evaluated long-context capabilities using the following tasks: NarrativeQA (Kočiskỳ et al., 2018), Scrolls-Qasper, Scrolls-Quality (Shaham et al., 2022), and XLSum (Hasan et al., 2021). Additionally, PG-19 (Rae et al., 2019) and Isabelle (Wu et al., 2022) are another common evaluation tasks among long-context Transformer papers (Rae et al., 2019; Wu et al., 2022; Chen et al., 2023). Detailed information on the evaluation data is provided in Table 10 in Section D.2.
Figure 4: Normalized scores of long-context tasks evaluated by Flan 1B with different Transformer models and different memory segment configurations. FAM outperforms all other BSWA configurations.
We evaluated the long-context capabilities of the 1B TransformerBSWA model trained in Section 3.1 using memory segment sizes ranging from 0 to 8. As shown in Figure 4, TransformerFAM outperformed TransformerBSWA on all the long context tasks (LCT), regardless of the number of memory segments in BSWA. It shows a significant performance improvement on ScrollsQasper and NarrativeQA, where it has to understand 5k to 500k tokens of context before answering a question. The LCT results demonstrate that TransformerFAM can effectively compress and retain important contextual information within extremely long contexts.
Above M1, the number of memory segments does not significantly impact LCT performance on TransformerBSWA, because the input sequences are much longer than the window size of all experiments. We observed the same phenomenon in TransformerFAM, and TransformerFAM uses $3$ memory segments in Figure 4. The figure shows the normalized scores of all tasks to view the scores on the same scale. The raw results are in Table 11 in Section D.2.
We further evaluated TransformerFAM and TransformerBSWA on 8B and 24B models. As shown in Table 1, TransformerFAM demonstrates scalability as the model size increases. This suggests that self-attention can route local information relevant to each input sequence while simultaneously routing contextual information to the FAM.
@@
In addition, TransformerFAM marginally surpasses TransformerBSWA on GPT-3 tasks (Brown et al., 2020) (see Table 2). This is not expected because all tasks are shorter than 2k tokens. We hypothesize this improvement stems from the efficient contextual representation by TransformerFAM. By offloading contextual data to FAM, TransformerFAM reduces duplication within input activations, optimizing latent space usage.
@@
Table 2: GPT-3 Task Report: Summarizes GPT-3 performance on ranking and generative tasks. (Details in Table 12)
Thus, BSWA memory segments (local representation) and FAM (global representation) complement each other. For LLMs, we recommend using FAM for compressed contextual representation alongside BSWA memory segments up to inference budgets (e.g., 2k, 8k, 32k (Team et al., 2023), or 128k (Achiam et al., 2023)).
3.4 Ablation study
3.4.1 FAM length
In the Flan 1B model, we observed performance saturation on ScrollsQasper, ScrollsQuality, and XLSum tasks when the FAM length reached 64. Interestingly, performance declined when FAM length exceeded 64. This suggests that information compression is more effective when space is limited.
FAM Len. ScrollsQasper ScrollsQuality XLSum
4 5.0 27.1 15.1
16 6.0 25.2 15.2
64 7.2 27.9 15.9
256 5.1 26.5 16.0
1024 5.3 26.3 16.0
Table 3: LCT scores according to FAM Length
3.4.2 The number of previous FAM blocks
In Figure 2(b), the input query attends to the FAM as denoted Attention to Feedback. The input query can attend to not only the immediately previous FAM, but also to more previous FAMs. Table 4 shows the XLSum scores for different numbers of previous FAM blocks. As the table shows, increasing the number of blocks did not have a significant effect, because the previous FAM already encapsulates all the previous information by a feedback copy. Therefore, the default setup attends to only the immediately previous FAM.
FAM blocks 1 2 3 4 6
XLSum 15.9 15.8 15.7 15.1 16.2
Table 4: XLSum scores according to the number of FAM blocks
In Section B.4, additional ablation studies are presented, due to the page limitation.
@@ 4 Related Work
There have been attempts to incorporate feedback mechanisms into the Transformer, but most of them involve feeding the output activations from the top layer to the bottom (Bulatov et al., 2022; Chevalier et al., 2023) or to intermediate layers (Fan et al., 2020). Since the top three layers in the Transformer are heavily focused on output reconstruction (Pasad et al., 2021), we hypothesize that there is a significant representational gap between the top and other layers. In this paper, we propose a feedback mechanism between intermediate layers.
There were papers that compressed information blockwise (Rae et al., 2019; Guo et al., 2019; Gupta & Berant, 2020; Mohtashami & Jaggi, 2023; Mu et al., 2023). However, in those papers, the information was not propagated infinitely. Relevant prior work includes the use of recurrent cross-attention between blocks (Hutchins et al., 2022), enabling the propagation of compressed information to subsequent blocks. Additionally, incorporating feedback from a few upper layers has been used to integrate past information (Ju et al., 2022). We propose TransformerFAM under the assumption that the human brain processes homogenous, heterogeneous, and feedback data with the same attention mechanism across distributed brain areas.
In Appendix E, additional related works are presented, due to the page limitation.
@@ 5 Conclusion
In the film ’Memento’ (2000), the protagonist struggles with anterograde amnesia, which means he can not remember anything before happened in the last 10 minutes, but his long-term memory is intact, He has to tattoo important information on his body to remember it. This is similar to the current state of large language models (LLMs). LLMs memorize the entire internet thanks to scaling laws (Kaplan et al., 2020), which allow them to store an enormous amount of information in large weights (long-term memory). However, their short-term memory is limited by the attention window. As a result, the complex prompt engineering becomes necessary to help them recall important details. We propose a new architecture called TransformerFAM that could fix anterograde amnesia of LLMs.
@@ A.1 Flan-PaLM architecture
Table 5 provides detailed information on the architecture of Flan-PaLM for 1B, 8B, and 24B models. MQA (Multi-Query Attention) (Shazeer, 2019) is an attention mechanism that employs a single set of keys and values for all attention heads.
@@ B.1 Initial FAM
When FAM are used as queries, they summarize the block context. Therefore, we learn by adding learnable summarization embeddings to the token embedding lookup level of the Transformer model. This is the same as prepending learnable embeddings in soft prompt tuning (Lester et al., 2021). The difference is that full attention is applied between the FAM prompt activations, and the updated FAM is used for the next block. The FAM prompt is passed to the next transformer layer through the forward propagation of the Transformer model, and it has a summary representation that is suitable for the level of the layer.
Prefix tuning (Li & Liang, 2021) can also be used to train learnable initial FAM at each Transformer layer. However, we had difficulty matching the magnitude of the learnable prefix to the regular input sequence, and the results of prompt tuning were consistently better. Ablation study in Section B.4.4 shows prompt tuning outperforms. In addition, prefix tuning has the disadvantage of adding additional weights of FAM length to each layer.
@@ B.2.1 FAM Position Encoding
We used rotary position embedding (RoPE) (Su et al., 2024) in all of our experiments. Each input sequence is assigned an integer position $m$, which is converted into sinusoidal encoding in the form of $\exp(im\theta)$ where $\theta=10000^{-2i/d_{model}}$ (Vaswani et al., 2017).
@@ B.2.2 Random Position Offset
The input length extrapolation problem of Transformer is well known (Chen et al., 2023; Xiong et al., 2023). For example, a Transformer LLM trained with 2k tokens experiences a severe performance drop when generating 8k tokens. This is because machine learning (ML) does not generalize well to situations that it has not seen during training.
@@ B.2.3 Random State Passing
The same problem existed in RNNs, and Random State Passing (RSP) (Narayanan et al., 2019) was proposed to handle long-form speech in the speech domain. We also used RSP to generalize the number of updates of FAM.
@@
On the other hand, saved FAM can be thought of as learnable prefix for prefix tuning (Li & Liang, 2021). It is also possible to train only the FAM while the model is frozen and use them for various downstream tasks or personalization. This part is left as a future research topic.
@@ B.3 FAM illustrated
In Section 2.2, we define the feedback loop as feedback activations that feed contextual representation back into each block of BSWA. This feedback loop is formally described in Algorithm 2. While Figure 2(b) illustrates Algorithm 2, the static image makes it challenging to fully grasp the dynamic nature of the decoding self-attention mechanism. To clarify this, we create a multi-frame animation in Figure 5 that demonstrates how the attention mechanism evolves over time (top to bottom).
@@ B.4 Ablation Studies
B.4.1 FAM Position Encoding
@@
Figure 6 shows FAM (i.e. Last number in Table 7) outperforming FAM-POS (i.e. Zero) in PG-19 accuracy over most base frequencies.
B.4.2 Random Position Offset
As mentioned in Section B.2.2, the input length extrapolation problem of the Transformer is well-known. The ”Attention is All You Need” (Vaswani et al., 2017) introduced sinusoidal position encoding in the form of $\exp(i(m-n)\theta)$ where $\theta=10000^{-2i/d_{model}}$. Popular solutions for full attention models include increasing the base frequency from 10k to 500k (Chen et al., 2023) or scaling down $(m-n)$ to $(m-n)\times(1024/4096)$ (Xiong et al., 2023).
@@ B.4.2 Random Position Offset
In Figure 6, FAM shows better PG-19 accuracy than FAM-RPO at the 10k base frequency. As mentioned in Section D.2, the max length of PG-19 was truncated to 256k in the experiments.
@@
In addition, we did not observe the attention sink phenomenon (Xiao et al., 2023) in our TransformerBSWA experiments. The paper proposes that Transformers with sliding window attention should store initial tokens in a KV cache for long-context inference. However, the relative position of SWA is restricted to a range of 0 to window size, independent of input sequence length. Our TransformerBSWA implementation, trained on 8.5k tokens, successfully operated for inference up to 256k tokens without any issues.
B.4.3 Random State Passing
To extrapolate the number of FAM updates, Section B.2.3 proposes Random State Passing (RSP). In Figure 6, FAM shows significantly better PG-19 accuracy than FAM-RSP. This demonstrates that RSP plays a crucial role in training FAM.
@@ B.4.3 Random State Passing
B.4.4 Prompt vs. Prefix
Section B.1 mentions that Prompt tuning outperforms Prefix tuning for training FAM in terms of performance, memory, and number of parameters. Figure 6 shows that Prefix tuning for FAM training leads to performance degradation in PG-19 (FAM vs. FAM+Prefix).
B.4.5 TransformerXL vs. TransformerBSWA
@@
B.4.6 Comparison with other methods
We compared FAM with Recurrent Memory Transformer (RMT) (Bulatov et al., 2022), because RMT also implements memory in Transformer using feedback mechanism from top layer to bottom layer. As shown in Figure 8, RMT showed worse performance than Block Sliding Window Attention (BSWA) with 1 memory segment, in the PassKey retrieval task. FAM solved the PassKey retrieval task, but RMT did not work at all with very long filler token lengths.
@@ B.4.6 Comparison with other methods
AutoCompressors (Chevalier et al., 2023) is an extension of RMT that continuously accumulates blockwise memory of RMT. AutoCompressors theoretically should be able to solve the PassKey task since it maintains all the memory tokens for all blocks. However, as shown in Figure 8, its performance drops sharply after 18k tokens. This is because the model only saw up to 18k tokens during training. It fails to generalize to longer filler token lengths. The AutoCompressors in Figure 8 accumulates 260 memories to support up to 260k tokens.

### v3

@@ 3.2 PassKey Retrieval
The PassKey retrieval task is a recent benchmark used in several long-context transformer papers [71, 68, 67]. In this task, a passkey is presented at the beginning, followed by a very long filler context. Finally, a question about the passkey is asked, as shown in Fig. 10 in Section D.1.
@@
As shown in Fig. 3(a), TransformerFAM was able to perfectly solve the task with a filler context of up to 260k tokens. In the figure, MX denotes the number of BSWA memory segments. The performance of TransformerBSWA improves significantly up to M2, after which it saturates. The performance of M12 also drops significantly after 20k tokens. The theoretical receptive field of M2 is 36k (i.e. depth(18) $\times$ memory segment(2k)), but the effective receptive field is much shorter.
(a) PassKey Retrieval
(b) Long Context Tasks
Figure 3: (a) PassKey Retrieval: Performance across different Transformer models and memory segment configurations. MX denotes the number of BSWA memory segments. FAM represents TransformerFAM with 0 memory segments. TransformerFAM successfully solves the task. (b) LCT: Normalized scores of long-context tasks evaluated by Flan 1B with different Transformer models and different memory segment configurations. FAM outperforms all other BSWA configurations.
In Fig. 3(a), it is important to compare M1_SG and M1. M1_SG has a stop gradient applied to one memory segment, which limits the receptive field to the window size. This is because the model cannot learn which contextual information stored in the memory segment will have a good result later. SWA with back propagation through time (BPTT) functions similarly to a time-limited RNN. It is common to use stop gradients on the memory segments of TransformerXL [26, 63]. However, we recommend against this practice.
In Section B.4.8, we compare our work with recent Transformer with memory papers [57, 63].
@@ 3.3 Long Context Tasks
Gemini [69] evaluated long-context capabilities using the following tasks: NarrativeQA [19], Scrolls-Qasper, Scrolls-Quality [60], and XLSum [51]. Additionally, PG-19 [21] and Isabelle [56] are another common evaluation tasks among long-context Transformer papers [21, 56, 67]. Detailed information on the evaluation data is provided in Table 10 in Section D.2.
We evaluated the long-context capabilities of the 1B TransformerBSWA model trained in Section 3.1 using memory segment sizes ranging from 0 to 8. As shown in Fig. 3(b), TransformerFAM outperformed TransformerBSWA on all the long context tasks (LCT), regardless of the number of memory segments in BSWA. It shows a significant performance improvement on ScrollsQasper and NarrativeQA, where it has to understand 5k to 500k tokens of context before answering a question. The LCT results demonstrate that TransformerFAM can effectively compress and retain important contextual information within extremely long contexts.
Above M1, the number of memory segments does not significantly impact LCT performance on TransformerBSWA, because the input sequences are much longer than the window size of all experiments. We observed the same phenomenon in TransformerFAM, and TransformerFAM uses $3$ memory segments in Fig. 3(b). The figure shows the normalized scores of all tasks to view the scores on the same scale. The raw results are in Table 11 in Section D.2.
We further evaluated TransformerFAM and TransformerBSWA on 8B and 24B models. As shown in Table 1, TransformerFAM demonstrates scalability as the model size increases. This suggests that self-attention can route local information relevant to each input sequence while simultaneously routing contextual information to the FAM. However, the performance improvements are not substantial, indicating room for further enhancements in working memory mechanisms.
@@
In addition, TransformerFAM marginally surpasses TransformerBSWA on GPT-3 tasks [33] (see Table 2). This result is unexpected since all tasks involve sequences shorter than 2k tokens. We hypothesize that this improvement arises from the efficient contextual representation by TransformerFAM. By offloading contextual data to FAM, TransformerFAM reduces redundancy within input activations, optimizing latent space usage. [^1] It can be viewed as the decoder analog of register tokens [78] in ViT encoders [32] to process global context.
@@
Table 2: Summarizes GPT-3 performance on ranking and generative tasks. (Details in Table 12)
Thus, BSWA memory segments (local representation) and FAM (global representation) complement each other. For LLMs, we recommend using FAM for compressed contextual representation alongside BSWA memory segments up to inference budgets (e.g., 2k, 8k, 32k [69], or 128k [70]). Due to page limitations in the main paper, ablation studies are presented in Section B.4.
@@ 4 Related Work
There have been attempts to incorporate feedback mechanisms into the Transformer, but most of them involve feeding the output activations from the top layer to the bottom [57, 63] or to intermediate layers [27]. Since the top three layers in the Transformer are heavily focused on output reconstruction [45], we hypothesize that there is a significant representational gap between the top and other layers. In this paper, we propose a feedback mechanism between intermediate layers.
There were papers that compressed information blockwise [21, 25, 31, 71, 72]. However, in those papers, the information was not propagated infinitely. Relevant prior work includes the use of recurrent cross-attention between blocks [52], enabling the propagation of compressed information to subsequent blocks. Additionally, incorporating feedback from a few upper layers has been used to integrate past information [61]. We propose TransformerFAM under the assumption that the human brain processes homogenous, heterogeneous, and feedback data with the same attention mechanism across distributed brain areas. Additional related works are presented in Appendix E.
@@ 5 Conclusion
In the film ’Memento’ (2000), the protagonist struggles with anterograde amnesia, which means he can not remember anything before happened in the last 10 minutes, but his long-term memory is intact, He has to tattoo important information on his body to remember it. This is similar to the current state of large language models (LLMs). LLMs memorize the entire internet thanks to scaling laws [39], which allow them to store an enormous amount of information in large weights (long-term memory). However, their short-term memory is limited by the attention window. As a result, the complex prompt engineering becomes necessary to help them recall important details. We propose a new architecture called TransformerFAM that could fix anterograde amnesia of LLMs.
@@ A.1 Flan-PaLM architecture
Table 3 provides detailed information on the architecture of Flan-PaLM for 1B, 8B, and 24B models. MQA (Multi-Query Attention) [22] is an attention mechanism that employs a single set of keys and values for all attention heads.
@@ B.1 FAM initialization
When FAM are used as queries, they summarize the block context. Therefore, we learn by adding learnable summarization embeddings to the token embedding lookup level of the Transformer model. This is the same as prepending learnable embeddings in soft prompt tuning [49]. The difference is that full attention is applied between the FAM prompt activations, and the updated FAM is used for the next block. The FAM prompt is passed to the next transformer layer through the forward propagation of the Transformer model, and it has a summary representation that is suitable for the level of the layer.
Prefix tuning [50] can also be used to train learnable initial FAM at each Transformer layer. However, we had difficulty matching the magnitude of the learnable prefix to the regular input sequence, and the results of prompt tuning were consistently better. Ablation study in Section B.4.6 shows prompt tuning outperforms. In addition, prefix tuning has the disadvantage of adding additional weights of FAM length to each layer.
@@ B.2.1 FAM Position Encoding
We used rotary position embedding (RoPE) [80] in all of our experiments. Each input sequence is assigned an integer position $m$, which is converted into sinusoidal encoding in the form of $\exp(im\theta)$ where $\theta=10000^{-2i/d_{model}}$ [12].
@@ B.2.2 Random Position Offset
The input length extrapolation problem of Transformer is well known [67, 65]. For example, a Transformer LLM trained with 2k tokens experiences a severe performance drop when generating 8k tokens. This is because machine learning (ML) does not generalize well to situations that it has not seen during training.
@@ B.2.3 Random State Passing
The same problem existed in RNNs, and Random State Passing (RSP) [24] was proposed to handle long-form speech in the speech domain. We also used RSP to generalize the number of updates of FAM.
@@
On the other hand, saved FAM can be thought of as learnable prefix for prefix tuning [50]. It is also possible to train only the FAM while the model is frozen and use them for various downstream tasks or personalization. This part is left as a future research topic.
@@ B.3 FAM illustrated
In Section 2.2, we define the feedback loop as feedback activations that feed contextual representation back into each block of BSWA. This feedback loop is formally described in Algorithm 2. While Fig. 2(b) illustrates Algorithm 2, the static image makes it challenging to fully grasp the dynamic nature of the decoding self-attention mechanism. To clarify this, we create a multi-frame animation in Fig. 4 that demonstrates how the attention mechanism evolves over time (top to bottom).
@@ B.4 Ablation Studies
B.4.1 FAM length
In the Flan 1B model, we observed performance saturation on ScrollsQasper, ScrollsQuality, and XLSum tasks when the FAM length reached 64. Interestingly, performance declined when FAM length exceeded 64, suggesting that information compression is more effective with limited space. This constraint on memory capacity is reminiscent of Miller’s Law [1], which posits that the average person can only hold approximately $7(\pm 2)$ items in their working memory at any given time.
FAM Len. ScrollsQasper ScrollsQuality XLSum
4 5.0 27.1 15.1
16 6.0 25.2 15.2
64 7.2 27.9 15.9
256 5.1 26.5 16.0
1024 5.3 26.3 16.0
Table 5: LCT scores according to FAM Length
B.4.2 The number of previous FAM blocks
In Fig. 2(b), the input query attends to the FAM as denoted Attention to Feedback. The input query can attend to not only the immediately previous FAM, but also to more previous FAMs. Table 6 shows the XLSum scores for different numbers of previous FAM blocks. As the table shows, increasing the number of blocks did not have a significant effect, because the previous FAM already encapsulates all the previous information by a feedback copy. Therefore, the default setup attends to only the immediately previous FAM.
FAM blocks 1 2 3 4 6
XLSum 15.9 15.8 15.7 15.1 16.2
Table 6: XLSum scores according to the number of FAM blocks
B.4.3 FAM Position Encoding
@@
Fig. 5 shows FAM (i.e. Last number in Table 7) outperforming FAM-POS (i.e. Zero) in PG-19 accuracy over most base frequencies.
B.4.4 Random Position Offset
As mentioned in Section B.2.2, the input length extrapolation problem of the Transformer is well-known. The "Attention is All You Need" [12] introduced sinusoidal position encoding in the form of $\exp(i(m-n)\theta)$ where $\theta=10000^{-2i/d_{model}}$. Popular solutions for full attention models include increasing the base frequency from 10k to 500k [67] or scaling down $(m-n)$ to $(m-n)\times(1024/4096)$ [65].
@@ B.4.4 Random Position Offset
In Fig. 5, FAM shows better PG-19 accuracy than FAM-RPO at the 10k base frequency. As mentioned in Section D.2, the max length of PG-19 was truncated to 256k in the experiments.
@@
In addition, we did not observe the attention sink phenomenon [66] in our TransformerBSWA experiments. The paper proposes that Transformers with sliding window attention should store initial tokens in a KV cache for long-context inference. However, the relative position of SWA is restricted to a range of 0 to window size, independent of input sequence length. Our TransformerBSWA implementation, trained on 8.5k tokens, successfully operated for inference up to 256k tokens without any issues.
B.4.5 Random State Passing
To extrapolate the number of FAM updates, Section B.2.3 proposes Random State Passing (RSP). In Fig. 5, FAM shows significantly better PG-19 accuracy than FAM-RSP. This demonstrates that RSP plays a crucial role in training FAM.
@@ B.4.5 Random State Passing
B.4.6 Prompt vs. Prefix
Section B.1 mentions that Prompt tuning outperforms Prefix tuning for training FAM in terms of performance, memory, and number of parameters. Fig. 5 shows that Prefix tuning for FAM training leads to performance degradation in PG-19 (FAM vs. FAM+Prefix).
B.4.7 TransformerXL vs. TransformerBSWA
@@
B.4.8 Comparison with other methods
We compared FAM with Recurrent Memory Transformer (RMT) [57], because RMT also implements memory in Transformer using feedback mechanism from top layer to bottom layer. As shown in Fig. 7, RMT showed worse performance than Block Sliding Window Attention (BSWA) with 1 memory segment, in the PassKey retrieval task. FAM solved the PassKey retrieval task, but RMT did not work at all with very long filler token lengths.
@@ B.4.8 Comparison with other methods
AutoCompressors [63] is an extension of RMT that continuously accumulates blockwise memory of RMT. AutoCompressors theoretically should be able to solve the PassKey task since it maintains all the memory tokens for all blocks. However, as shown in Fig. 7, its performance drops sharply after 18k tokens. This is because the model only saw up to 18k tokens during training. It fails to generalize to longer filler token lengths. The AutoCompressors in Fig. 7 accumulates 260 memories to support up to 260k tokens.